Tim Bleckie is the owner of of Beckie Investment and Real-estate company The company recently purchased four tracts of land in holy farms estate and six tracts in newburg woods The tracts are all equally desirable and sell for about the same amount.
a. What is the probability that the next two tracts will be sold in Newburg Woods?
b. What is the probability that of the next four sold at least one will be in Holly Farms?
c. Are these events independent or dependent?

Answer by stanbon(75887)

Tim Bleckie is the owner of of Beckie Investment and Real-estate company The company recently purchased four tracts of land in holy farms estate and six tracts in newburg woods The tracts are all equally desirable and sell for about the same amount.
a. What is the probability that the next two tracts will be sold in Newburg Woods?
---
Ans: 6C2/10C2 = 1/3
-----------------------------------
b. What is the probability that of the next four sold at least one will be in Holly Farms?
Ans: 1 - P(none are from Holly Farms) = 1 - [6C4/10C4] = 1 - 0.0714 = 0.9286
-----------------------------------
c. Are these events independent or dependent?
Procedure:
If P(a and b) = P(a)*P(b) then a and b are independent
If P(a|b) = P(a)*P(b) then a and b are independent.
-----------------------------------
Cheers,
Stan H.